OpenStudy (anonymous):

How do you come to the conclusion that when considering a equilateral triangle inside a circumscribed circle, the height of the triangle is 2/3 of the radius of the circle?

OpenStudy (dumbcow):

 the height is actually greater than the radius sin(30) = h/r = 1/2 --> h = r/2 therefore the height of triangle is 3/2 the radius